OpenStudy (lucaz):

with initial vel. 700 m/s, what is the angle for a launch to reach 12,500 m? (horizontally)

OpenStudy (aaronq):

Just take into account the component of the vector in the x direction.  \(v_{i~x}=cos\theta v_i\)

OpenStudy (lucaz):

yeah, but dont we need the time?

OpenStudy (lucaz):

time taken to travel the given distance?

OpenStudy (aaronq):

hm yeah, you're right. Maybe it's possible to use the vertical velocity, solve for time, and plug that into the x-motion equation?

OpenStudy (lucaz):

I tried to find t when v(y) is zero, but then what is the angle... I dont know

OpenStudy (aaronq):

\(\sf v_f=v_i+a\Delta t\) \(\sf 0=sin\theta v_i-g\Delta t\rightarrow \Delta t=\dfrac{sin\theta v_i}{g}\) Then: \(\sf\Delta x=cos\theta v_i\Delta t +\dfrac{1}{2}(0)(\Delta t)^2\) \(\sf\Delta x=cos\theta v_i(\dfrac{sin\theta v_i}{g})=\dfrac{cos\theta sin \theta v_i}{g} \) but then you'd have to use some trig

OpenStudy (aaronq):

that's what i was thinking, but i'm not entirely sure that assuming that the final velocity is 0 in the y direction is sound.

OpenStudy (aaronq):

I forgot to square the velocity btw

OpenStudy (lucaz):

my book has this formula with sin and cos.. I will try again, thank you

OpenStudy (aaronq):

no problem, good luck!